OpenStudy (anonymous):

Let an integer be chosen at random from the integers 1 to 30 inclusive. Find the probability that the integer chosen is divisible by 4 or is greater than 25 and less than 30. 1/3 11/30 2/5

OpenStudy (anonymous):

@yordicecream

OpenStudy (freckles):

P(A or B)=P(A)+P(B)-P(A and B)

OpenStudy (anonymous):

1/3?

OpenStudy (freckles):

well how many multiples of 4 are there between 1 and 30?

OpenStudy (anonymous):

7

OpenStudy (freckles):

ok and how many integers are between 25 and 30

OpenStudy (anonymous):

4

OpenStudy (freckles):

ok and how many of those numbers that were a multiple of 40 that is also between 25 and 30

OpenStudy (freckles):

multiple of 30*

OpenStudy (freckles):

\[P(A \text{ or } B)=P(A)+P(B)-P(A \text{ and } B)=\frac{7}{30}+\frac{4}{30}-\frac{?}{30}\]

OpenStudy (freckles):

we just need that last number there

OpenStudy (freckles):

4(1),4(2),4(3),4(4),4(5),4(6),4(7) lol I said multiple of 30

OpenStudy (freckles):

how many of those numbers are between 25 an 30

OpenStudy (anonymous):

30?

OpenStudy (freckles):

4(1)=4 4(2)=8 4(3)=12 4(4)=16 4(7)=28

OpenStudy (freckles):

hmmm the only number i see between 25 and 30 is...

OpenStudy (freckles):

28

OpenStudy (freckles):

that is 1 number

OpenStudy (freckles):

\[P(A \text{ or } B)=P(A)+P(B)-P(A \text{ and } B)=\frac{7}{30}+\frac{4}{30}-\frac{1}{30}\]

OpenStudy (anonymous):

10/30

OpenStudy (anonymous):

but theres no 10/30 in the options... so i think its 1/3 since its equivalent to 10/30

OpenStudy (freckles):

I think you know that 10/30 reduces to 1/3 in your deepest of hearts

OpenStudy (anonymous):

so im right?

OpenStudy (freckles):

divide top and bottom by 10 \[\frac{10}{30} \frac{\frac{1}{10}}{\frac{1}{10}}=\frac{1}{3}\]

OpenStudy (anonymous):

thnx